OpenStudy (anonymous):

A line passes through (2, –1) and (8, 4).Write an equation for the line in point-slope form. Rewrite the equation in standard form using integers

OpenStudy (kc_kennylau):

1. Find its slope 2. Express the line in point-slope form 3. Convert it to standard form

OpenStudy (anonymous):

y + 1 = 5/6 (x + 2); –5x + 6y = –16 y – 2 =5/6 (x + 1); –5x + 6y = 17 y + 1 =5/6 (x – 2); –5x + 6y = –16 y – 1 = 5/6 (x – 2); –5x + 6y = 16 what would be the best answer ??

OpenStudy (marigirl):

prior to finding the equation, let us fist f ind the gradient by using the formula (y2-y1/x2-x1) so the gradient will be m=5/6 now to obtain the equation let us use y-y1=m(x-x1) lets use the coordinate (2, –1) and plug in the information y-(-1)=5/6(x-2) now lets clean up...what do we get?

OpenStudy (anonymous):

wow thats alot of wok lol i was just hoping for a quick response

OpenStudy (marigirl):

well u can see from your multiple answer options that 5/6 is consistent throughout the options therefore, that is one less thing for u to figure out......its a matter of u putting in the numbers in the equation y-y1=m(x-x1)

OpenStudy (anonymous):

so what would be the right choice for the answer

OpenStudy (marigirl):

y + 1 =5/6 (x – 2)